# ../svg_images/ Quantum Architecture Search with Unsupervised Representation Learning

### v3

@@ Abstract
Unsupervised representation learning presents new opportunities for advancing Quantum Architecture Search (QAS) on Noisy Intermediate-Scale Quantum (NISQ) devices. QAS is designed to optimize quantum circuits for Variational Quantum Algorithms (VQAs). Most QAS algorithms tightly couple the search space and search algorithm, typically requiring the evaluation of numerous quantum circuits, resulting in high computational costs and limiting scalability to larger quantum circuits. Predictor-based QAS algorithms mitigate this issue by estimating circuit performance based on structure or embedding. However, these methods often demand time-intensive labeling to optimize gate parameters across many circuits, which is crucial for training accurate predictors. Inspired by the classical neural architecture search algorithm Arch2vec, we investigate the potential of unsupervised representation learning for QAS without relying on predictors. Our framework decouples unsupervised architecture representation learning from the search process, enabling the learned representations to be applied across various downstream tasks. Additionally, it integrates an improved quantum circuit graph encoding scheme, addressing the limitations of existing representations and enhancing search efficiency. This predictor-free approach removes the need for large labeled datasets. During the search, we employ REINFORCE and Bayesian Optimization to explore the latent representation space and compare their performance against baseline methods. Our results demonstrate that the framework efficiently identifies high-performing quantum circuits with fewer search iterations.
@@ 1 Introduction
Quantum Computing (QC) has made significant progress over the past decades. Advances in quantum hardware and new quantum algorithms have demonstrated potential advantages (Stein et al., 2023) over classical computers in various tasks, such as image processing (Wang et al., 2022), reinforcement learning (Skolik et al., 2022), knowledge graph embedding (Ma et al., 2019), and network architecture search (Zhang et al., 2022; Giovagnoli et al., 2023; Du et al., 2022). However, the scale of quantum computers is still limited by environmental noise, which leads to unstable performance. These noisy intermediate-scale quantum (NISQ) devices lack fault tolerance, which is not expected to be achieved in the near future (Preskill, 2018). The variational quantum algorithm (VQA), a hybrid quantum algorithm that utilizes quantum operations with adjustable parameters, is considered a leading strategy in the NISQ era (Cerezo et al., 2021). In VQA, the parameterized quantum circuit (PQC) with trainable parameters is viewed as a general paradigm of quantum neural networks and has achieved notable success in quantum machine learning. These parameters control quantum circuit operations, adjusting the distribution of circuit output states, and are updated by a classical optimizer based on a task-specific objective function. Although VQA faces challenges such as Barren Plateaus (BP) and scalability issues, it has demonstrated the potential to improve performance across various domains, including image processing, combinatorial optimization, chemistry, and physics (Pramanik et al., 2022; Amaro et al., 2022; Tilly et al., 2022). One example of a VQA is the variational quantum eigensolver (VQE) (Peruzzo et al., 2014; Tilly et al., 2022), which approximates the ground state and offers flexibility for quantum machine learning. We are considering using VQE to evaluate the performance of certain quantum circuits.
@@
Unsupervised representation learning seeks to discover hidden patterns or structures within unlabeled data, a well-studied problem in computer vision research (Radford et al., 2015). One common approach is the autoencoder, which is effective for feature representation. It consists of an encoder and decoder, which first maps images into a compact feature space and then decodes them to reconstruct similar images. Beyond images, autoencoders can also learn useful features from graphs, such as encoding and reconstructing directed acyclic graphs (DAGs) or neural network architectures (Yan et al., 2020; Zhang et al., 2019; Pan et al., 2018; Wang et al., 2016). In most research, architecture search and representation learning are coupled, which results in inefficient searches heavily dependent on labeled architectures that require numerous evaluations. The Arch2vec framework aims to decouple representation learning from architecture search, allowing downstream search algorithms to operate independently (Yan et al., 2020). This decoupling leads to a smooth latent space that benefits various search algorithms without requiring extensive labeling.
@@
Quantum architecture search (QAS) or quantum circuit architecture search is a framework for designing quantum circuits efficiently and automatically, aiming to optimize circuit performance (Du et al., 2022). Various algorithms have been proposed for QAS (Zhang et al., 2022; Du et al., 2022; Zhang et al., 2021; He et al., 2023a; Giovagnoli et al., 2023). However, most algorithms combine the search space and search algorithm, leading to inefficiency and high evaluation costs. The effectiveness of the search algorithm often depends on how well the search space is defined, embedded, and learned. Finding a suitable circuit typically requires evaluating different architectures many times. Although predictor-based QAS He et al. (2023a) can separate representation learning from the search algorithm, it often relies on labeling different architectures via evaluation, and the training performance depends heavily on the quantity and quality of evaluations and the embedding. In this work, we are inspired by the idea of decoupling, and we aim to conduct QAS without labeling. We seek to explore whether decoupling can embed quantum circuit architectures into a smooth latent space, benefiting predictor-free QAS algorithms.We summarise our contributions as follows:
@@
Extensive experiments on quantum machine learning tasks, including quantum state preparation, max-cut, and quantum chemistry (Liang et al., 2019; Poljak & Rendl, 1995; Tilly et al., 2022), confirm the effectiveness of our framework. The pre-trained quantum architecture embeddings significantly enhance QAS across these applications.
@@ 2 Related Work
Graph data is becoming a crucial tool for understanding complex interactions between real-world entities, such as biochemical molecules (Jiang et al., 2021), social networks (Shen et al., 2023), purchase networks from e-commerce platforms (Li et al., 2021), and academic collaboration networks (Newman, 2001). Graphs are typically represented as discrete data structures, making it challenging to solve downstream tasks due to large search spaces. Our work focuses on unsupervised graph representation learning, which seeks to embed graphs into low-dimensional, compact, and continuous representations without supervision while preserving the topological structure and node attributes. In this domain, approaches such as those proposed by Perozzi et al. (2014); Wang et al. (2016); Grover & Leskovec (2016); Tang et al. (2015) use local random walk statistics or matrix factorization-based objectives to learn graph representations. Alternatively, methods like Kipf & Welling (2016); Hamilton et al. (2017) reconstruct the graph’s adjacency matrix by predicting edge existence, while others, such as Veličković et al. (2018); Sun et al. (2019); Peng et al. (2020), maximize the mutual information between local node representations and pooled graph representations. Additionally, Xu et al. (2019) investigate the expressiveness of Graph Neural Networks (GNNs) in distinguishing between different graphs and introduce Graph Isomorphism Networks (GINs), which are shown to be as powerful as the Weisfeiler-Lehman test (Leman & Weisfeiler, 1968) for graph isomorphism. Inspired by the success of Arch2vec (Yan et al., 2020), which employs unsupervised graph representation learning for classical neural architecture search (NAS), we adopt GINs to injectively encode quantum architecture structures, as quantum circuit architectures can also be represented as DAGs.
@@
As discussed in the previous section, PQCs are essential as ansatz for various VQAs (Benedetti et al., 2019). The expressive power and entangling capacity of PQCs play a crucial role in their optimization performance (Sim et al., 2019). Poorly designed ansatz can suffer from limited expressive power or entangling capacity, making it difficult to reach the global minimum for an optimization problem. Moreover, such ansatz may be more prone to noise (Stilck França & Garcia-Patron, 2021), inefficiently utilize quantum resources, or lead to barren plateaus that hinder the optimization process (McClean et al., 2018; Wang et al., 2021). To address these challenges, QAS has been proposed as a systematic approach to identify optimal PQCs. The goal of QAS is to efficiently and effectively search for high-performance quantum circuits tailored to specific problems, minimizing the loss functions while adhering to constraints such as hardware qubit connections, native quantum gate sets, quantum noise models, training loss landscapes, and other practical considerations. Quantum architectures share many properties with neural network architectures, such as hierarchical, directed, and acyclic structures. As a result, QAS methods have been heavily inspired by techniques from NAS. Specifically, approaches such as greedy algorithms (Mitarai et al., 2018; Tang et al., 2021), evolutionary or genetic methods (Zhang & Zhao, 2022; Ding & Spector, 2022), RL-based engines (Kuo et al., 2021; Ostaszewski et al., 2021), Bayesian optimization (Duong et al., 2022), and gradient-based methods (Zhang et al., 2022) have all been employed to discover improved PQCs for VQAs. However, these methods require the evaluation of numerous quantum circuits, which is both time-consuming and computationally expensive. To mitigate this issue, predictor-based approaches (Zhang et al., 2021; He et al., 2023b) have been introduced, but they also face limitations. These approaches rely on large sets of labeled circuits to train predictors with generalized capabilities and introduce additional uncertainty into the search process, necessitating the reevaluation of candidate circuits. In this work, we propose a framework aimed at further addressing these challenges.
@@ 3.1 Circuit Encoding Scheme
We represent quantum circuits as DAGs using the circuit encoding scheme , as described in He et al. (2023b; a). Each circuit is transformed into a DAG by mapping the gates on each qubit to a sequence of nodes, with two additional nodes added to indicate the input and output of circuits. The resulting DAG is described by an adjacency matrix, as shown in Figure 1(a). The set of nodes is further characterized by a gate matrix, which shows the node features including position information.
@@ 3.2.1 Preliminaries
The most common graph autoencoders (GAEs) consist of an encoder and a decoder, where the encoder maps a graph into a feature space, and the decoder reconstructs the graph from those features. One prominent example is the variational graph autoencoder (VGAE), a promising framework for unsupervised graph representation learning that utilizes a graph convolutional network as its encoder and a simple inner product as its decoder (Kipf & Welling, 2016). In this work, however, we do not employ the common VGAE as a framework for learning latent representations. Instead, we utilize a more powerful encoder GIN (Xu et al., 2019).
@@ 3.3.1 Reinforcement Learning (RL)
After conducting initial trials with PPO (Schulman et al., 2017) and A2C (Huang et al., 2022), we adopt REINFORCE (Williams, 1992) as a more effective reinforcement learning algorithm for architecture search. In this approach, the environment’s state space consists of pre-trained embeddings, and the agent uses a one-cell LSTM as its policy network. The agent selects an action, corresponding to a sampled latent vector based on the distribution of the current state, and transitions to the next state based on the chosen action. The reward for max-cut and quantum chemistry tasks is defined as the ratio of energy to ground energy, with values outside the range [0, 1] clipped to 0 or 1. For the state preparation task, circuit fidelity is used as the reward. We employ an adaptive batch size, with the number of steps per training epoch determined by the average reward of the previous epoch. Additionally, we use a linear adaptive baseline, defined by the formula , where denotes the baseline, is a predefined value in the range [0,1], and is the average reward. Each run in this work involves 1000 searches.
@@ 3.3.2 Bayesian Optimization (BO)
As another search strategy used in this work without labeling, we employ Deep Networks for Global Optimization (DNGO)(Snoek et al., 2015) in the context of BO. We adopt a one-layer adaptive BO regression model with a basis function extracted from a feed-forward neural network, consisting of 128 units in the hidden layer, to model distributions over functions. Expected Improvement (EI)(Mockus, 1977) is selected as the acquisition function. EI identifies the top-k embeddings for each training epoch, with a default objective value of 0.9. The training begins with an initial set of 16 samples, and in each subsequent epoch, the top-k architectures proposed by EI are added to the batch. The network is retrained for 100 epochs using the architectures from the updated batch. This process is iterated until the predefined number of search iterations is reached.
@@ 4.1 Pre-training Performance
Observation (1): GAE and VGAE (Kipf & Welling, 2016) are two popular baselines for NAS. In an attempt to find models capable of capturing superior latent representations of quantum circuit architectures, we initially applied these two well-known models. However, due to the increased complexity of quantum circuit architectures compared to neural network architectures, these models failed to deliver the expected results. In contrast, models based on GINs (Xu et al., 2019) successfully obtained valid latent representations, attributed to their more effective neighbor aggregation scheme. Table 1 presents a performance comparison between the original model using the encoding and the improved model with our enhanced encoding for 4, 8, and 12 qubit circuits, evaluated across five metrics: Accuracy, which measures the reconstruction accuracy of gate types in the gate matrix for the held-out test set; Accuracy, which reflects the reconstruction accuracy of qubits that the gates act on; Accuracy, which measures the reconstruction accuracy of the adjacency matrix; Falpos, which represents the mean false positives in the reconstructed adjacency matrix; and KLD (KL divergence), which indicates the continuity and smoothness of the latent representation. The results in the table indicate that the improved model with our enhanced encoding achieves comparable or better than the original. This improvement can be attributed to two factors: first, the new encoding better captures the specific characteristics of the circuits, and second, the fusion of outputs from multiple layers of GIN helps retain shallow information, resulting in more stable training.
@@
Observation (2): In Figure 2, we employ two popular techniques, PCA (Shlens, 2014) and t-SNE (Van der Maaten & Hinton, 2008), to visualize the high-dimensional latent representations of 4- and 12-qubit quantum machine learning (QML) applications based on our pre-trained models. The results highlight the effectiveness of our new encoding approach for unsupervised clustering and high-dimensional data visualization. The figures show that the latent representation space of quantum circuits is smooth and compact, with architectures of similar performance clustering together when the search space is limited to 4 qubits. Notably, high-performance quantum circuit architectures are concentrated on the right side of the visualizations. In particular, PCA yields exceptionally smooth and compact representations with strong clustering effects, making it easier and more efficient to conduct QAS within such a structured latent space. This provides a robust foundation for our QAS algorithms.

### v4

# Review:Quantum Architecture Search with Unsupervised Representation Learning

@@ Abstract
Unsupervised representation learning presents new opportunities for advancing Quantum Architecture Search (QAS) on Noisy Intermediate-Scale Quantum (NISQ) devices. QAS is designed to optimize quantum circuits for Variational Quantum Algorithms (VQAs). Most QAS algorithms tightly couple the search space and search algorithm, typically requiring the evaluation of numerous quantum circuits, resulting in high computational costs and limiting scalability to larger quantum circuits. Predictor-based QAS algorithms mitigate this issue by estimating circuit performance based on structure or embedding. However, these methods often demand time-intensive labeling to optimize gate parameters across many circuits, which is crucial for training accurate predictors. Inspired by the classical neural architecture search algorithm Arch2vec, we investigate the potential of unsupervised representation learning for QAS without relying on predictors. Our framework decouples unsupervised architecture representation learning from the search process, enabling the learned representations to be applied across various downstream tasks. Additionally, it integrates an improved quantum circuit graph encoding scheme, addressing the limitations of existing representations and enhancing search efficiency. This predictor-free approach removes the need for large labeled datasets. During the search, we employ REINFORCE and Bayesian Optimization to explore the latent representation space and compare their performance against baseline methods. We further validate our approach by executing the best-discovered MaxCut circuits on IBM’s ibm_sherbrooke quantum processor, confirming that the architectures retain optimal performance even under real hardware noise. Our results demonstrate that the framework efficiently identifies high-performing quantum circuits with fewer search iterations.
Researchers and practitioners in quantum machine learning, quantum architecture search, and quantum circuit optimization, particularly those interested in applying unsupervised learning techniques to improve efficiency and scalability of circuit design for near-term quantum devices (NISQ).
@@ 1 Introduction
Quantum Computing (QC) has made significant progress over the past decades. Advances in quantum hardware and new quantum algorithms have demonstrated potential advantages [1] over classical computers in various tasks, such as image processing [2], reinforcement learning [3], knowledge graph embedding [4], and network architecture search [5, 6, 7]. However, the scale of quantum computers is still limited by environmental noise, which leads to unstable performance. These noisy intermediate-scale quantum (NISQ) devices lack fault tolerance, which is not expected to be achieved in the near future [8]. The variational quantum algorithm (VQA), a hybrid quantum algorithm that utilizes quantum operations with adjustable parameters, is considered a leading strategy in the NISQ era [9]. In VQA, the parameterized quantum circuit (PQC) with trainable parameters is viewed as a general paradigm of quantum neural networks and has achieved notable success in quantum machine learning. These parameters control quantum circuit operations, adjusting the distribution of circuit output states, and are updated by a classical optimizer based on a task-specific objective function. Although VQA faces challenges such as Barren Plateaus (BP) and scalability issues, it has demonstrated the potential to improve performance across various domains, including image processing, combinatorial optimization, chemistry, and physics [10, 11, 12]. One example of a VQA is the variational quantum eigensolver (VQE) [13, 12], which approximates the ground state and offers flexibility for quantum machine learning. We are considering using VQE to evaluate the performance of certain quantum circuits.
@@
Unsupervised representation learning seeks to discover hidden patterns or structures within unlabeled data, a well-studied problem in computer vision research [14]. One common approach is the autoencoder, which is effective for feature representation. It consists of an encoder and decoder, which first maps images into a compact feature space and then decodes them to reconstruct similar images. Beyond images, autoencoders can also learn useful features from graphs, such as encoding and reconstructing directed acyclic graphs (DAGs) or neural network architectures [15, 16, 17, 18]. In most research, architecture search and representation learning are coupled, which results in inefficient searches heavily dependent on labeled architectures that require numerous evaluations. The Arch2vec framework aims to decouple representation learning from architecture search, allowing downstream search algorithms to operate independently [15]. This decoupling leads to a smooth latent space that benefits various search algorithms without requiring extensive labeling.
@@
Quantum architecture search (QAS) or quantum circuit architecture search is a framework for designing quantum circuits efficiently and automatically, aiming to optimize circuit performance [7]. Various algorithms have been proposed for QAS [5, 7, 19, 20, 6]. However, most algorithms combine the search space and search algorithm, leading to inefficiency and high evaluation costs. The effectiveness of the search algorithm often depends on how well the search space is defined, embedded, and learned. Finding a suitable circuit typically requires evaluating different architectures many times. Although predictor-based QAS [20] can separate representation learning from the search algorithm, it often relies on labeling different architectures via evaluation, and the training performance depends heavily on the quantity and quality of evaluations and the embedding. In this work, we are inspired by the idea of decoupling, and we aim to conduct QAS without labeling. We seek to explore whether decoupling can embed quantum circuit architectures into a smooth latent space, benefiting predictor-free QAS algorithms.We summarise our contributions as follows:
@@
Extensive experiments on quantum machine learning tasks, including quantum state preparation, max-cut, and quantum chemistry [21, 22, 12], confirm the effectiveness of our framework on simulator and real quantum hardware. The pre-trained quantum architecture embeddings significantly enhance QAS across these applications.
@@ 2 Related Work
Graph data is becoming a crucial tool for understanding complex interactions between real-world entities, such as biochemical molecules [23], social networks [24], purchase networks from e-commerce platforms [25], and academic collaboration networks [26]. Graphs are typically represented as discrete data structures, making it challenging to solve downstream tasks due to large search spaces. Our work focuses on unsupervised graph representation learning, which seeks to embed graphs into low-dimensional, compact, and continuous representations without supervision while preserving the topological structure and node attributes. In this domain, approaches such as those proposed by [27, 18, 28, 29] use local random walk statistics or matrix factorization-based objectives to learn graph representations. Alternatively, methods like [30, 31] reconstruct the graph’s adjacency matrix by predicting edge existence, while others, such as [32, 33, 34], maximize the mutual information between local node representations and pooled graph representations. Additionally, [35] investigate the expressiveness of Graph Neural Networks (GNNs) in distinguishing between different graphs and introduce Graph Isomorphism Networks (GINs), which are shown to be as powerful as the Weisfeiler-Lehman test [36] for graph isomorphism. Inspired by the success of Arch2vec [15], which employs unsupervised graph representation learning for classical neural architecture search (NAS), we adopt GINs to injectively encode quantum architecture structures, as quantum circuit architectures can also be represented as DAGs.
@@
As discussed in the previous section, PQCs are essential as ansatz for various VQAs [37]. The expressive power and entangling capacity of PQCs play a crucial role in their optimization performance [38]. Poorly designed ansatz can suffer from limited expressive power or entangling capacity, making it difficult to reach the global minimum for an optimization problem. Moreover, such ansatz may be more prone to noise [39], inefficiently utilize quantum resources, or lead to barren plateaus that hinder the optimization process [40, 41]. To address these challenges, QAS has been proposed as a systematic approach to identify optimal PQCs. The goal of QAS is to efficiently and effectively search for high-performance quantum circuits tailored to specific problems, minimizing the loss functions while adhering to constraints such as hardware qubit connections, native quantum gate sets, quantum noise models, training loss landscapes, and other practical considerations. Quantum architectures share many properties with neural network architectures, such as hierarchical, directed, and acyclic structures. As a result, QAS methods have been heavily inspired by techniques from NAS. Specifically, approaches such as greedy algorithms [42, 43], evolutionary or genetic methods [44, 45], RL-based engines [46, 47], Bayesian optimization [48], and gradient-based methods [5] have all been employed to discover improved PQCs for VQAs. However, these methods require the evaluation of numerous quantum circuits, which is both time-consuming and computationally expensive. To mitigate this issue, predictor-based approaches [19, 49] have been introduced, but they also face limitations. These approaches rely on large sets of labeled circuits to train predictors with generalized capabilities and introduce additional uncertainty into the search process, necessitating the reevaluation of candidate circuits. In this work, we propose a framework aimed at further addressing these challenges.
@@ 3.1 Circuit Encoding Scheme
We represent quantum circuits as DAGs using the circuit encoding scheme , as described in [49, 20]. Each circuit is transformed into a DAG by mapping the gates on each qubit to a sequence of nodes, with two additional nodes added to indicate the input and output of circuits. The resulting DAG is described by an adjacency matrix, as shown in Figure 1(a). The set of nodes is further characterized by a gate matrix, which shows the node features including position information.
@@ 3.2.1 Preliminaries
The most common graph autoencoders (GAEs) consist of an encoder and a decoder, where the encoder maps a graph into a feature space, and the decoder reconstructs the graph from those features. One prominent example is the variational graph autoencoder (VGAE), a promising framework for unsupervised graph representation learning that utilizes a graph convolutional network as its encoder and a simple inner product as its decoder [30]. In this work, however, we do not employ the common VGAE as a framework for learning latent representations. Instead, we utilize a more powerful encoder GIN [35].
@@ 3.3.1 Reinforcement Learning (RL)
After conducting initial trials with PPO [50] and A2C [51], we adopt REINFORCE [52] as a more effective reinforcement learning algorithm for architecture search. In this approach, the environment’s state space consists of pre-trained embeddings, and the agent uses a one-cell LSTM as its policy network. The agent selects an action, corresponding to a sampled latent vector based on the distribution of the current state, and transitions to the next state based on the chosen action. The reward for max-cut and quantum chemistry tasks is defined as the ratio of energy to ground energy, with values outside the range [0, 1] clipped to 0 or 1. For the state preparation task, circuit fidelity is used as the reward. We employ an adaptive batch size, with the number of steps per training epoch determined by the average reward of the previous epoch. Additionally, we use a linear adaptive baseline, defined by the formula , where denotes the baseline, is a predefined value in the range [0,1], and is the average reward. Each run in this work involves 1000 searches.
@@ 3.3.2 Bayesian Optimization (BO)
As another search strategy used in this work without labeling, we employ Deep Networks for Global Optimization (DNGO)[53] in the context of BO. We adopt a one-layer adaptive BO regression model with a basis function extracted from a feed-forward neural network, consisting of 128 units in the hidden layer, to model distributions over functions. Expected Improvement (EI)[54] is selected as the acquisition function. EI identifies the top-k embeddings for each training epoch, with a default objective value of 0.9. The training begins with an initial set of 16 samples, and in each subsequent epoch, the top-k architectures proposed by EI are added to the batch. The network is retrained for 100 epochs using the architectures from the updated batch. This process is iterated until the predefined number of search iterations is reached.
@@ 4.1 Pre-training Performance
Observation (1): GAE and VGAE [30] are two popular baselines for NAS. In an attempt to find models capable of capturing superior latent representations of quantum circuit architectures, we initially applied these two well-known models. However, due to the increased complexity of quantum circuit architectures compared to neural network architectures, these models failed to deliver the expected results. In contrast, models based on GINs [35] successfully obtained valid latent representations, attributed to their more effective neighbor aggregation scheme. Table 1 presents a performance comparison between the original model using the encoding and the improved model with our enhanced encoding for 4, 8, and 12 qubit circuits, evaluated across five metrics: Accuracy, which measures the reconstruction accuracy of gate types in the gate matrix for the held-out test set; Accuracy, which reflects the reconstruction accuracy of qubits that the gates act on; Accuracy, which measures the reconstruction accuracy of the adjacency matrix; Falpos, which represents the mean false positives in the reconstructed adjacency matrix; and KLD (KL divergence), which indicates the continuity and smoothness of the latent representation. The results in the table indicate that the improved model with our enhanced encoding achieves comparable or better than the original. This improvement can be attributed to two factors: first, the new encoding better captures the specific characteristics of the circuits, and second, the fusion of outputs from multiple layers of GIN helps retain shallow information, resulting in more stable training.
@@
Observation (2): In Figure 2, we employ two popular techniques, PCA [55] and t-SNE [56], to visualize the high-dimensional latent representations of 4- and 12-qubit quantum machine learning (QML) applications based on our pre-trained models. The results highlight the effectiveness of our new encoding approach for unsupervised clustering and high-dimensional data visualization. The figures show that the latent representation space of quantum circuits is smooth and compact, with architectures of similar performance clustering together when the search space is limited to 4 qubits. Notably, high-performance quantum circuit architectures are concentrated on the right side of the visualizations. In particular, PCA yields exceptionally smooth and compact representations with strong clustering effects, making it easier and more efficient to conduct QAS within such a structured latent space. This provides a robust foundation for our QAS algorithms.